OpenStudy (heart_offire):

in triangle ABC, A=(3x + 7), B =(2x-8), c=(5-x) solve for x

563blackghost (563blackghost):

If these are the angles of the triangle you would add them up which would eventually equal to 180. \(\huge\bf{(3x+7)+(2x-8)+(5-x)=180}\)

563blackghost (563blackghost):

\(\Large\bf{(3x+7)+(2x-8)+(5-x)=180}\) Sorry the other was to big

OpenStudy (thefatdog):

3x+7+2x-8+5-x=180 3x+2x-x+7-8+5=180 you can do the rest! :D

KjSaif (kjsaif):

Simplify both sides of the equation. (3x + 2x + −x) + (7 + − 8 + 5)=180(Combine Like Terms) 4x + 4=180

KjSaif (kjsaif):

Subtract 4 from both sides.

KjSaif (kjsaif):

4x + 4 − 4 = 180−4

KjSaif (kjsaif):

4x=176

KjSaif (kjsaif):

Divide both sides by 4.

KjSaif (kjsaif):

4x/4 =176/4

KjSaif (kjsaif):

x=44

KjSaif (kjsaif):

So.....Your answer is x = 44

OpenStudy (heart_offire):

OK thank you!, you were very helpful